# On the order of lazy cellular automata

### v2

@@ 3 Lazy CA with a quasi-constant active transition
Let $\tau:A^{G}\to A^{G}$ be a lazy cellular automaton with minimal local map $\mu:A^{S}\to A$, unique active transition $p\in A^{S}$, and writing symbol $a:=\mu(p)\in A\setminus\{p(e)\}$. In this section, we assume that $p$ is quasi-constant, which means that $p$ is not constant and there exists $r\in S$ such that $p|_{S\setminus\{r\}}$ is constant. We call $r\in S$ the non-constant element of $p$. Without loss of generality, we assume that $p\in A^{S}$ is defined as follows: for all $s\in S$,
$p(s):=\begin{cases}1&\text{ if }s=r\\ 0&\text{ if }s\neq r.\end{cases}$
Because of Corollary 3, we only need to consider the case when $a=p(s)$ for some $s\in S$ (as otherwise, $S_{a}=\emptyset$).
3.1 Non-constant element $r\neq e$
Proposition 3.
For any $n\geq 2$, $\textup{ord}\left(\tau\right)>n$ if and only if $r^{j}\notin S$ for all $2\leq j\leq n$.
@@ 3.1 Non-constant element $r\neq e$
Assume that $\textup{ord}\left(\tau\right)>n$. By Theorem 2, there exists a word $w=(s_{1},\ldots,s_{n-1})\in(S_{1})^{n-1}$ such that for all $i\leq j$,
$(s_{j}\cdots s_{i})^{-1}\notin S_{0}^{-1}S_{1}.$
By the construction of $p$ in this section, we have $S_{0}=S\setminus\{r\}$ and $S_{1}=\{r\}$, so the word $w$ has the form $w=(r,r,\dots r)$. Hence, the above condition is equivalent to $r^{j}\notin S$, for all $2\leq j\leq n$.
Conversely, assume that $r^{j}\notin S$ for all $2\leq j\leq n$. Define $x\in A^{G}$ as follows:
Claim 1.
For all $m\in\{0,\ldots,n-1\}$, $(g\cdot\tau^{m}(x))|_{S}=p$ if and only if $g=r^{-m}$.
Proof.
We prove this by strong induction on $m$. For $m=0$, it is clear by the construction of $x$. Now, assume the existence of $m\in\{0,\ldots,n-2\}$ such that the following holds
$(g\cdot\tau^{\ell}(x))|_{S}=p\iff g=r^{-\ell},\quad\forall\ell\in\{0,\ldots,m\}.$
We will show that $(g\cdot\tau^{m+1})(x)|_{S}=p$ if and only if $g=r^{-(m+1)}$. First we will show that $(r^{-(m+1)}\cdot\tau^{m+1})(x)|_{S}=p$. For any $s\in S$, (3) implies that
If $sr^{-(m+1)}=r^{-\ell}$, then $s=r^{m+1-\ell}$. Note that $m+1-\ell\in\{1,\ldots,m+2\}$, where $m+2\leq n$ by construction of $m$. Since $r^{j}\notin S$ for all $2\leq j\leq n$, it follows that $s=r^{m+1-\ell}$ if and only if $m+1-\ell=1$. As a consequence, $\ell=m$. Then, by (3) and (4)
$\tau^{m+1}(x)(sr^{-(m+1)})=1$ $\iff s=r,$
$\tau^{m+1}(x)(sr^{-(m+1)})=0$ $\iff s\in S\setminus\{r\}.$
Therefore, $(r^{-(m+1)}\cdot\tau^{m+1}(x))|_{S}=p$. Conversely, suppose that $(g\cdot\tau^{m+1}(x))|_{S}=p$ for some $g\in G$. It follows by construction of $p$ that
$\tau^{m+1}(x)(g)$ $=p(e)=0,$
$\tau^{m+1}(x)(rg)$ $=p(r)=1.$
It follows by the previous claim and Lemma 1 that $\tau^{j-1}\neq\tau^{j}$ for all $j\in\{1,\ldots,n\}$. Therefore, $\textup{ord}\left(\tau\right)>n$ by Corollary 2. ∎
Theorem 1 (2.) follows as a direct consequence of Proposition 3.
@@
When the universe $G$ has an element of infinite order, such as when $G=\mathbb{Z}^{d}$, the previous corollary implies that there is always a lazy CA of every given order $n\geq 2$.
@@ 3.2 Non-constant element $r=e$
For any $n\geq 2$, $\textup{ord}\left(\tau\right)>n$ if and only if there exists a word $(s_{1},\dots,s_{n-1})\in(S\setminus\{e\})^{n-1}$ such that
$(s_{j}\dots s_{i})^{-1}\not\in S,\quad\forall i\leq j.$
@@
Suppose that $\textup{ord}\left(\tau\right)>n$. By Theorem 2, there exists a word $w=(s_{1},\ldots,s_{n-1})\in(S_{0})^{n-1}$ such that for all $i\leq j$,
$(s_{j}\cdots s_{i})^{-1}\notin S_{1}^{-1}S_{0}.$
By the construction of $p$ in this section, we have that $S_{0}=S\setminus\{e\}$ and $S_{1}=\{e\}$, so the direct implication follows.
Conversely, assume that there exists a word $(s_{1},\dots,s_{n-1})\in(S\setminus\{e\})^{n-1}$ such that
$(s_{j}\dots s_{i})^{-1}\not\in S,\quad\forall i\leq j.$
Define $x\in A^{G}$ as follows: for all $g\in G$
@@
For all $i\in\{0,\ldots,n-1\}$, $(g\cdot\tau^{i}(x))|_{S}=p$ if and only if $g=s_{n-(i+1)}\cdots s_{1}s_{0}$.
@@
First of all, observe that the hypothesis $(s_{j}\dots s_{i})^{-1}\not\in S$, for all $i\leq j$ is equivalent to
$\{s_{n-j}\cdots s_{1}s_{0}\}_{j=k}^{n}\cap\bigcup_{i=1}^{k}\big((S\setminus\{e\})s_{n-i}\cdots s_{1}s_{0}\big)=\emptyset,\quad\forall 1\leq k\leq n.$ (5)
We prove Claim 2 by strong induction on $i$. For the base case $i=0$, fix $k=1$ in (5), so we see that $x(ss_{n-1}\cdots s_{1})=0$ for all $s\in S\setminus\{e\}$ and $x(s_{n-1}\cdots s_{1})=1$ by the construction of $x$. It follows that $(s_{n-1}\cdots s_{1}\cdot x)|_{S}=p$. Conversely, assume that $(g\cdot x)|_{S}=p$ for some $g\in G$, so $x(g)=1$ and $x(sg)=0$ for all $s\in S\setminus\{e\}$. If follows by the construction of $p$ that $g=s_{\ell}\cdots s_{1}s_{0}$ for some $\ell\in\{0,\ldots,n-1\}$ and $s_{j}\cdots s_{1}s_{0}\notin(S\setminus\{e\})g$ for all $j\in\{0,\ldots,n-1\}$. If $\ell<n-1$,
$s_{\ell+1}=(s_{\ell+1}\cdots s_{1}s_{0})(s_{\ell}\cdots s_{1}s_{0})^{-1}\notin S\setminus\{e\},$
@@
Now, assume the existence of $m\in\{0,\ldots,n-2\}$ such that the following holds
$(g\cdot\tau^{i}(x))|_{S}=p\iff g=s_{n-(i+1)}\cdots s_{1}s_{0},\quad\forall i\in\{0,\ldots,m\}.$
@@
$s_{j}\cdots s_{1}s_{0}\notin(S\setminus\{e\})s_{n-(m+2)}\cdots s_{1}s_{0},\quad\forall j=0,\ldots,n-(m+2).$
@@
Conversely, assume that $(g\cdot\tau^{m+1})(x)|_{S}=p$ for some $g\in G$. Observe that $\tau^{m+1}(x)(g)=1$ and $\tau^{m+1}(x)(sg)=0$ for all $s\in S\setminus\{e\}$. By (6), we have that $g=s_{\ell}\cdots s_{1}s_{0}$ for some $\ell\in\{0,\ldots,n-(m+2)\}$ and $s_{j}\cdots s_{1}s_{0}\notin(S\setminus\{e\})g$ for all $j\in\{0,\ldots,n-(m+2)\}$. If $\ell<n-(m+2)$, then
$s_{\ell+1}=(s_{\ell+1}\cdots s_{1}s_{0})(s_{\ell}\cdots s_{1}s_{0})^{-1}\notin S\setminus\{e\},$
@@
The previous claim and Corollary 1 imply that $\tau^{j-1}\neq\tau^{j}$ for all $j\in\{1,\ldots,n\}$. The result follows by Corollary 2. ∎

### v3

@@ 3 Lazy CA with a quasi-constant active transition
Recall that a pattern $p\in A^{S}$ is quasi-constant if it is not constant and there exists an element $r\in S$, called the non-constant element of $p$, such that $p|_{S\setminus\{r\}}$ is constant.
Theorem 2.
Let $\tau:A^{G}\to A^{G}$ be a lazy cellular automaton with unique active transition $p\in A^{S}$ and writing symbol $a\in A\setminus\{p(e)\}$. Assume that $p$ is quasi-constant with non-constant element $r\in S$.
  1. 1.
    If $a\neq p(s)$ for all $s\in S$, then $\textup{ord}\left(\tau\right)=2$.
  2. 2.
    $\textup{ord}\left(\tau\right)=\min\{n\geq 2:r^{n}\in S\}.$
  3. 3.
    If $r=e$ and $a=p(s)$ for all $s\in S\setminus\{e\}$, then $\textup{ord}\left(\tau\right)$ is finite if and only if there exists $n\geq 2$ such that for all words $w\in(S^{-1}\setminus\{e\})^{n-1}$ there exists a subword $v\sqsubseteq w$ such that $\theta(v)\in S$. In such case, the order of $\tau$ is the minimum $n$ satisfying this property.
Before proving this theorem, we shall mention some of its consequences. The following corollary corresponds to Theorem 2 in [3].
Corollary 4.
With the assumptions of Theorem 2, $\tau:A^{G}\to A^{G}$ is idempotent if and only if one of the following holds:
  1. 1.
    $a\neq p(s)$ for all $s\in S$.
  2. 2.
    $r\neq e$ and $r^{2}\in S$.
  3. 3.
    $r=e$ and $S=S^{-1}$.
@@
Parts (1.) and (2.) are clear consequences of Theorem 2. Observe that part (3.) in Theorem 2 with $n=2$ says that all $s\in S\setminus\{e\}$ satisfy $s^{-1}\in S\setminus\{e\}$, which is equivalent to $S=S^{-1}$. ∎
@@
Example 8.
Consider the the lazy CA $\tau:A^{\mathbb{Z}}\to A^{\mathbb{Z}}$ with minimal neighborhood $S=\{0,1,n\}$, with $n\geq 2$, unique active transition $p=01\_\dots\_n\in A^{S}$ and writing symbol $a=1\in A$. By Theorem 2 (2.), $\textup{ord}\left(\tau\right)=n$.
Example 9.
Consider the the lazy CA $\tau:A^{\mathbb{Z}}\to A^{\mathbb{Z}}$ with minimal neighborhood $S=\{-k,0,1,k\}$, with $k\geq 2$, unique active transition $p=0\_\dots\_10\_\dots\_0\in A^{S}$ and writing symbol $a=0\in A$. We want to find the smallest $n$ such that every word $w\in(-S\setminus\{0\})^{n-1}=\{-k,-1,k\}^{n-1}$ has a subword $v\sqsubseteq w$ such that $\theta(v)\in S$. If $k$ or k-k appear in the word $w$, then we may always take $v=k$ or v=kv=-k, respectively. Hence, assume that $w=(-1,\dots,-1)$. It is clear that we need the length of $w$ to be at least $k$ so we can find $v\sqsubseteq w$ such that $\theta(v)=-k\in S$. Therefore, $\textup{ord}\left(\tau\right)=k+1$.
In order to prove Theorem 2, let $\tau:A^{G}\to A^{G}$ be a lazy cellular automaton with minimal local map $\mu:A^{S}\to A$, unique active transition $p\in A^{S}$ that is quasi-constant with non-constant element $r\in S$, and writing symbol $a:=\mu(p)\in A\setminus\{p(e)\}$. Without loss of generality, we assume that $p\in A^{S}$ is defined as follows: for all $s\in S$,
$p(s):=\begin{cases}1&\text{ if }s=r\\ 0&\text{ if }s\neq r.\end{cases}$
Because of Corollary 3, we only need to consider the case when $a=p(s)$ for some $s\in S$ (as otherwise, $S_{a}=\emptyset$). The proof of Theorem 2 is divided in two cases: when the non-constant element $r$ is different from the group identity $e$, and when they are equal. The following sections deal with each of these cases.
3.1 Non-constant element $r\neq e$
Proposition 3.
For any $n\geq 2$, $\textup{ord}\left(\tau\right)>n$ if and only if $r^{j}\notin S$ for all $j\in[2,n]$.
Proof.
Assume that $\textup{ord}\left(\tau\right)>n$. By Theorem 1, there exists a word $w=(s^{-1}_{1},\ldots,s^{-1}_{n-1})\in(S_{1}^{-1})^{n-1}$ such that for all $i\leq j$,
$s_{i}^{-1}\dots s_{j}^{-1}\notin S_{0}^{-1}S_{1}.$
By the construction of $p$ in this section, we have $S_{0}=S\setminus\{r\}$ and $S_{1}=\{r\}$, so the word $w$ has the form $w=(r^{-1},r^{-1},\dots,r^{-1})$. Hence, the above condition is equivalent to
$r^{-k}\notin(S^{-1}\setminus\{r\})r,\quad\forall k\in[1,n-1],\quad\Longleftrightarrow\quad r^{j}\notin S,\quad\forall j\in[2,n].$
Conversely, assume that $r^{j}\notin S$ for all $j\in[2,n]$. Define $x\in A^{G}$ as follows:
Claim 1.
For all $m\in[0,n-1]$, $(g\cdot\tau^{m}(x))|_{S}=p$ if and only if $g=r^{-m}$.
Proof.
We prove this by strong induction on $m$. For $m=0$, it is clear by the construction of $x$. Now, assume the existence of $m\in[0,n-2]$ such that
$(g\cdot\tau^{\ell}(x))|_{S}=p\iff g=r^{-\ell},\quad\forall\ell\in[0,m].$
We will show that $(g\cdot\tau^{m+1})(x)|_{S}=p$ if and only if $g=r^{-(m+1)}$. First we show that $(r^{-(m+1)}\cdot\tau^{m+1})(x)|_{S}=p$. For any $s\in S$, (4) implies that
If $sr^{-(m+1)}=r^{-\ell}$, then $s=r^{m+1-\ell}$. Note that $m+1-\ell\in[1,m+2]$, where $m+2\leq n$ by the choice of $m\in[0,n-2]$. Since $r^{j}\notin S$ for all $j\in[2,n]$, it follows that $s=r^{m+1-\ell}$ if and only if $m+1-\ell=1$. As a consequence, $\ell=m$. Then, by (4) and (5)
$\tau^{m+1}(x)(sr^{-(m+1)})=1$ $\iff s=r,$
$\tau^{m+1}(x)(sr^{-(m+1)})=0$ $\iff s\in S\setminus\{r\}.$
Therefore, $(r^{-(m+1)}\cdot\tau^{m+1}(x))|_{S}=p$. Conversely, suppose that $(g\cdot\tau^{m+1}(x))|_{S}=p$ for some $g\in G$. It follows by the construction of $p$ that
$\tau^{m+1}(x)(g)$ $=p(e)=0,$
$\tau^{m+1}(x)(rg)$ $=p(r)=1.$
It follows by the previous claim and Lemma 1 that $\tau^{n-1}\neq\tau^{n}$. Therefore, $\textup{ord}\left(\tau\right)>n$ by Corollary 2 and Proposition 2. ∎
Theorem 2 (2.) follows as a direct consequence of Proposition 3.
@@ 3.2 Non-constant element $r=e$
For any $n\geq 2$, $\textup{ord}\left(\tau\right)>n$ if and only if there exists a word $(s^{-1}_{1},\dots,s^{-1}_{n-1})\in(S^{-1}\setminus\{e\})^{n-1}$ such that
$s_{i}^{-1}\dots s_{j}^{-1}\not\in S,\quad\forall i\leq j.$
@@
Suppose that $\textup{ord}\left(\tau\right)>n$. By Theorem 1, there exists a word $w=(s^{-1}_{1},\dots,s^{-1}_{n-1})\in(S_{0}^{-1})^{n-1}$ such that for all $i\leq j$,
$s_{i}^{-1}\dots s_{j}^{-1}\notin S_{1}^{-1}S_{0}.$
By the construction of $p$ in this section, we have that $S_{0}=S\setminus\{e\}$ and $S_{1}=\{e\}$, so $s_{i}^{-1}\dots s_{j}^{-1}\notin S\setminus\{e\}$ for all $i\leq j$. As $s^{-1}_{i}\in S^{-1}\setminus\{e\}$, for all $i$, it is not possible that $s_{i}^{-1}=e$. If $s_{i}^{-1}\dots s_{j}^{-1}=e$ for some $i<j$, then $s_{i+1}^{-1}\dots s_{j}^{-1}=s_{i}\in S\setminus\{e\}$ which is a contradiction. Therefore, $s_{i}^{-1}\dots s_{j}^{-1}\notin S$ for all $i\leq j$, and the direct implication follows.
Conversely, assume that there exists a word $(s^{-1}_{1},\dots,s^{-1}_{n-1})\in(S^{-1}\setminus\{e\})^{n-1}$ such that
$s_{i}^{-1}\dots s_{j}^{-1}\not\in S,\quad\forall i\leq j.$
Define $x\in A^{G}$ as follows: for all $g\in G$,
@@
For all $i\in[0,n-1]$, $(g\cdot\tau^{i}(x))|_{S}=p$ if and only if $g=s_{n-(i+1)}\cdots s_{1}s_{0}$.
@@
First of all, observe that the hypothesis $s_{i}^{-1}\dots s_{j}^{-1}\not\in S$, for all $i\leq j$ implies that
$\{s_{n-j}\cdots s_{1}s_{0}\}_{j=k}^{n}\cap\bigcup_{i=1}^{k}\big((S\setminus\{e\})s_{n-i}\cdots s_{1}s_{0}\big)=\emptyset,\quad\forall k\in[1,n].$ (6)
Indeed, if $s_{n-j}\cdots s_{1}s_{0}\in(S\setminus\{e\})s_{n-i}\cdots s_{1}s_{0}$, then $s_{n-j+1}^{-1}\dots s_{n-i}^{-1}\in S\setminus\{e\}$, contradicting the hypothesis.
We prove Claim 2 by strong induction on $i$. For the base case $i=0$, fix $k=1$ in (6), so we see that $x(ss_{n-1}\cdots s_{1})=0$ for all $s\in S\setminus\{e\}$ and $x(s_{n-1}\cdots s_{1})=1$ by the construction of $x$. It follows that $(s_{n-1}\cdots s_{1}\cdot x)|_{S}=p$. Conversely, assume that $(g\cdot x)|_{S}=p$ for some $g\in G$, so $x(g)=1$ and $x(sg)=0$ for all $s\in S\setminus\{e\}$. If follows by the construction of $p$ that $g=s_{\ell}\cdots s_{1}s_{0}$ for some $\ell\in[0,n-1]$ and $s_{j}\cdots s_{1}s_{0}\notin(S\setminus\{e\})g$ for all $j\in[0,n-1]$. If $\ell<n-1$,
$s_{\ell+1}=(s_{\ell+1}\cdots s_{1}s_{0})(s_{0}^{-1}s_{1}^{-1}\dots s_{\ell}^{-1})\notin S\setminus\{e\},$
@@
For the induction hypothesis, assume the existence of $m\in[0,n-2]$ such that the following holds
$(g\cdot\tau^{i}(x))|_{S}=p\iff g=s_{n-(i+1)}\cdots s_{1}s_{0},\quad\forall i\in[0,m].$
@@
$s_{j}\cdots s_{1}s_{0}\notin(S\setminus\{e\})s_{n-(m+2)}\cdots s_{1}s_{0},\quad\forall j\in[0,n-(m+2)].$
@@
Conversely, assume that $(g\cdot\tau^{m+1})(x)|_{S}=p$ for some $g\in G$. Observe that $\tau^{m+1}(x)(g)=1$ and $\tau^{m+1}(x)(sg)=0$ for all $s\in S\setminus\{e\}$. By (7), we have that $g=s_{\ell}\cdots s_{1}s_{0}$ for some $\ell\in[0,n-(m+2)]$ and $s_{j}\cdots s_{1}s_{0}\notin(S\setminus\{e\})g$ for all $j\in[0,n-(m+2)]$. If $\ell<n-(m+2)$, then
$s_{\ell+1}=(s_{\ell+1}\cdots s_{1}s_{0})(s_{0}^{-1}s_{1}^{-1}\dots s_{\ell}^{-1})\notin S\setminus\{e\},$
@@
The previous claim and Corollary 1 imply that $\tau^{n-1}\neq\tau^{n}$. The result follows by Corollary 2 and Proposition 2. ∎